# PEARL: Self-Evolving Assistant for Time Management with Reinforcement Learning

### v1

@@ 7 Conclusion
In this work, we study calendar conflict resolution, a long-horizon, preference-driven decision-making task. We introduce CalConflictBench for systemic investigation, and evaluation results shows that current LLM agents degrade as horizons grow and conflicts become denser. To address this, We propose PEARL, a RL framework with explicit memory module and round-wise rewards, achieving strong gains on CalConflictBench.
@@
Our study is an initial step toward systematically evaluating and training preference-evolving agents for calendar conflict resolution, and it leaves several limitations for future work. First, CalConflictBench represents user preferences via structured, role-conditioned rules over event attributes, which makes evaluation reproducible but inevitably incomplete. In real-world settings, decisions can be driven by transient and hard-to-observe factors that are not reflected in calendar metadata—e.g., “I’m not in the mood for meetings today,” fatigue, stress, interpersonal dynamics, or unexpected urgent tasks. Such affective and situational signals are difficult to simulate faithfully and may only be expressed through natural language messages or behavioral cues. Consequently, agents that perform well in our benchmark may still fail under implicit, rapidly shifting drivers of user choices. Second, while we conduct all the necessary experiments to support our main claims, computational and time constraints prevent an exhaustive sweep over all possible combinations of evaluation parameters. Third, because current LLMs have limited context windows, we only evaluate histories of up to 20 past events. We leave designing principled mechanisms for dynamically selecting and summarizing relevant context over long horizons as future work.

### v2

@@ 7 Conclusion
In this work, we study calendar conflict resolution, a long-horizon, preference-driven decision-making task. We introduce CalConflictBench for systematic investigation, and evaluation results show that current LLM agents degrade as horizons grow and conflicts become denser. To address this, we propose PEARL, a RL framework with an explicit memory module and round-wise rewards, achieving strong gains on CalConflictBench.
@@
Our study is an initial step toward systematically evaluating and training preference-evolving agents for calendar conflict resolution, and it leaves several limitations for the future work. First, CalConflictBench represents user preferences via structured, role-conditioned rules over event attributes, which makes evaluation reproducible but inevitably incomplete. In real-world settings, decisions can be driven by transient and hard-to-observe factors that are not reflected in calendar metadata—e.g., “I’m not in the mood for meetings today,” fatigue, stress, interpersonal dynamics, or unexpected urgent tasks. Such affective and situational signals are difficult to simulate faithfully and may only be expressed through natural language messages or behavioral cues. Consequently, agents that perform well in our benchmark may still fail under implicit, rapidly shifting drivers of user choices. Second, while we conduct all the necessary experiments to support our main claims, computational and time constraints prevent an exhaustive sweep over all possible combinations of evaluation parameters. Third, because current LLMs have limited context windows, we only evaluate histories of up to 20 past events. We leave designing principled mechanisms for dynamically selecting and summarizing relevant context over long horizons as future work.